# AI-based Framework for Robust Model-Based Connector Mating in Robotic Wire Harness Installation

### v1

@@ III AI-BASED FRAMEWORK FOR AUTOMATED CONNECTOR MATING
Figure 2: AI-based Framework for Visuotactile Connector Mating: An initial robot program is created by the programmer using industry-standard tools ARTM (left). During the ramp-up phase, the robot executes the program repeatedly with varying search parameterizations. The resulting dataset is used to train MuTT a predictive visuotactile model of the robot and environment dynamics (center). This model then serves as a predictor for the first-order optimizer SPI which generates parameters that optimize the program for robust and fast connector mating given the observed process variance and the current environment (right).
@@ IV EXPERIMENTS
  • $L$ = Trajectory length, $F$ = Force Z (N)
@@ V RESULTS
  • $L$ = Trajectory length, $P$ = Position (mm), $F$ = Force Z (N)
@@ V-A Prediction Accuracy
The shadow program is trained in a supervised manner for 100 epochs using an initial learning rate of $5\text{x}10^{-5}$ with linear decay. Table II presents the prediction performance of MuTT across the five connectors. MuTT accurately predicts the expected skill execution for all five connectors, achieving an average trajectory position deviation of less than 1 mm and a force deviation within 1 N, which is on the order of sensor noise, despite the very diverse training dataset. Furthermore, MuTT predicts the expected search success with an F1 score exceeding $99~{}\%$ and demonstrates robust learning of the minority class (search failure) in the imbalanced datasets. The predicted trajectory length deviates by 20 points from the real world trajectory on average, corresponding to a deviation by one second from ground-truth cycle times. Overall prediction accuracies are sufficient for use in model-based optimization (Sec. V-B).
@@ V-C Ablation Study
  • $L$ = Trajectory length, $P$ = Position (mm), $F$ = Force Z (N)
  • CT = Cycle time (s), SR = Success rate (%)
[…]

### v2

@@ III AI-BASED FRAMEWORK FOR AUTOMATED CONNECTOR MATING
Figure 2: AI-based Framework for Visuotactile Connector Mating: A programmer creates a initial robot program using industry-standard tools (ARTM, left). During the ramp-up phase, the robot executes the program repeatedly with varying search parameterizations. The resulting dataset is used to train MuTT, a predictive visuotactile model of the robot and environment dynamics (center). MuTT serves as a sim2real predictor for the first-order optimizer SPI, which optimizes program parameters for robust and fast connector mating given the observed process variance and current environment image (right).
@@ IV EXPERIMENTS
  • Dataset metrics: number of training/test samples (# train/test), successful/failed trials (Succ/Fail), and mean ($\mu$) and std. dev. ($\sigma$) of trajectory length ($L$, in points) and end-effector Z-force ($F$, in N).
@@ V RESULTS
  • Evaluation of predicted task success, trajectory length ($L$, points), Cartesian position ($P$, mm), and end-effector Z-force ($F$, N).
@@ V-A Prediction Accuracy
The shadow program is trained in a supervised manner for 100 epochs using an initial learning rate of $5\text{x}10^{-5}$ with linear decay. Table II presents the prediction performance of MuTT across the five connectors. MuTT accurately predicts the expected skill execution for all five connectors, achieving an mean absolute error from the real trajectory (MAE) of the end-effector position of less than 1 mm and within 1 N of the end-effector force, which is on the order of FT sensor noise. Furthermore, MuTT predicts the expected search success with an F1 score exceeding $99~{}\%$ and demonstrates robust learning of the minority class (search failure) in the imbalanced datasets. The predicted trajectory length $L$ deviates by 20 points from the real world trajectory on average, corresponding to a deviation by one second from ground-truth cycle times. Overall prediction accuracies are sufficient for use in model-based optimization (Sec. V-B).
@@ V-C Ablation Study
  • See Table II and III for notation.